OpenStudy (anonymous):

find the 2 consecutive integers for 131

OpenStudy (anonymous):

this is what i got

OpenStudy (anonymous):

66,65

OpenStudy (anonymous):

n+(n+1)=131 2n+1=131 1-1=131-1 = 130 2n/2=130/2 n=65 sooo...65,66

OpenStudy (anonymous):

is it correct

OpenStudy (anonymous):

Yes it is 66,65

OpenStudy (anonymous):

yay thank you im going to post another one of these up to make sure im right on a different question